cos^2y+sinxcosxcosy p=sinycos^2x using u=sinx and v=siny
1
Expert's answer
2020-10-26T19:43:32-0400

We can rewrite the equality as:

"(1-v^2)\pm u\sqrt{1-u^2}\sqrt{1-v^2}p=v(1-u^2)"

The sign is chosen depending on the sign of expression "cos\,x\,cos\,y" .